OpenStudy (anonymous):

[(2x^3 + 3x^2 +28) / ((x+2)(x^2))] integrate 4rm 1 to 2

OpenStudy (anonymous):

please help me?

OpenStudy (anonymous):

with steps if you dont mind.

OpenStudy (anonymous):

not as bad as it seems. divide, \[\frac{2x^3+3x^2+28}{x^3+2x^2}\] is what you need

OpenStudy (anonymous):

ca you divide it with the steps. shown.

OpenStudy (anonymous):

i will try but it is hard to write division here. let me write it first with paper

OpenStudy (anonymous):

ah ok.

OpenStudy (anonymous):

well actually you get a quotient of 2, and a remainder of \[-x^2+28\]

OpenStudy (anonymous):

2 _______________ x^3 + 2x^2 | 2x^3+3x^2 + 28 2x^3 + 4x^2 -x^2 + 28

OpenStudy (anonymous):

that is the best way i can write it here

OpenStudy (anonymous):

just like long division, you say x^3 goes in to 2x^3 twice, put the 2 up top, multiply out and subtract

OpenStudy (anonymous):

that means v multiply because it says integrate from 1 to 2 is it?

OpenStudy (anonymous):

meaning \[\frac{2x^3+3x^2+28}{x^3+2x^2}=2+\frac{-x^2+28}{x^3+2x^2}\]

OpenStudy (anonymous):

we did not get anywhere close to integration yet, but yes, when you integrate 2 from 1 to 2 you will get 2 we still have to find the antiderivative of the second part, and for this we need "partial fractions"

OpenStudy (anonymous):

\[\frac{-x^2+28}{x^2(x+2)}=\frac{a}{x}+\frac{b}{x^2}+\frac{c}{x+2}\] and now we need a, b and c

OpenStudy (anonymous):

do you know how to do the partial fraction decomposition?

OpenStudy (anonymous):

nope i am new to this, what is the answer finally?

OpenStudy (anonymous):

for the partial fraction you will get \[\frac{-x^2+28}{x^2(x+2)}=\frac{14}{x}-\frac{7}{x^2}+\frac{6}{x+2}\]

OpenStudy (anonymous):

so in total \[\frac{2x^3+3x^2+28}{x^3+2x^2}=2+\frac{14}{x}-\frac{7}{x^2}+\frac{6}{x+2}\]

OpenStudy (anonymous):

now you have to integrate term by term to find the anti derivative

OpenStudy (anonymous):

you get \[\int 2+\frac{14}{x}-\frac{7}{x^2}+\frac{6}{x+2} dx = 2x+14\ln(x)+\frac{7}{x}+6\ln(x+2)\]

OpenStudy (anonymous):

thats is not the hard part, that you just take the anti derivative term by term. the hard part is finding that \[\frac{-x^2+28}{x^2(x+2)}=\frac{14}{x}-\frac{7}{x^2}+\frac{6}{x+2}\]

OpenStudy (anonymous):

hm ok bro thanks..